OpenStudy (highschoolmom2010):

What is the product in simplest form? State any restrictions on the variable.

OpenStudy (highschoolmom2010):

\[\frac{ y^2 }{ y-3 }*\frac{ y^2-y-6 }{ y^2+y}\]

OpenStudy (anonymous):

factor and then cancel out what you can. for instance y^2 -y - 6 =(y+2) (1-3)

OpenStudy (anonymous):

Factor everything first that can be factored

OpenStudy (highschoolmom2010):

 like that